# Stepwise Alignment for Constrained Language Model Policy Optimization

### v1

@@ 4 Stepwise Alignment for Constrained Policy Optimization
We propose an algorithm called Stepwise Alignment for Constrained Policy Optimization (SACPO) to solve the constrained policy optimization problem (10). SACPO takes a stepwise approach for reward and safety alignments of LLMs; that is, an LLM is first aligned for reward and then for safety (or vice versa). This operation is backed by theory, the resulting optimal LM policy is guaranteed to be identical to the one aligned with reward and safety metrics simultaneously. In addition, by taking the stepwise approach, we can enjoy several practical advantages of enhanced flexibility regarding algorithms and datasets. SACPO is outlined in Algorithm 1. Though we formulate a problem with a single safety function for simplicity, the arguments in this section are valid in the case of multiple safety functions. For more details, see Appendix B.
SACPO is based on Lagrangian methods [6] and uses a standard Lagrangian
$L(\pi,\lambda,\beta)\coloneqq R(\pi,\beta)+\lambda(G(\pi)-b),$ (11)
where $\pi\in\Pi$ is the primal variable and $\lambda\in\mathbb{R}_{+}$ is a dual variable or the Lagrangian multiplier. Note that, for any dual variable $\lambda\in\mathbb{R}_{+}$, by using a composite function $r^{\star}+\lambda g^{\star}$, we can convert the original constrained policy optimization problem (10) into the following max-min problem:
$\max_{\pi}\min_{\lambda\geq 0}\ L(\pi,\lambda,\beta)=\mathbb{E}_{\rho,\pi}\bigl{[}r^{\star}(x,y)+\lambda g^{\star}(x,y)\bigr{]}-\beta\mathbb{D}_{\mathrm{KL}}[\,\pi_{\theta}(y\mid x)\,\|\,\pi_{\mathrm{ref}}(y\mid x)\,]-\lambda b.$
@@
Practically, it is not hard to obtain such a conservative policy $\overline{\pi}$. If the usefulness (i.e., reward $r$) can be ignored, it is easy to acquire policies that refuse to generate potentially unsafe answers and output safe answers conservatively. Based on Assumption 1, we recall the following two lemmas regarding the strong duality and bounded optimal dual variable.
@@
For the proofs, see Appendix D. Our problem setting is a special case of those in typical constrained Markov decision process (CMDP, [2]) literature. Thus, Lemma 1 follows from Theorem 3 in Paternain et al. [30], and Lemma 2 follows from Lemma 1 in Ding et al. [13].
@@ 4.1 Optimal Policy Can be Directly Obtained from Reward-aligned Policy
3:  Choose the loss function $\mathcal{L}_{r}$ within $\{\mathcal{L}_{\text{DPO}},\mathcal{L}_{\text{IPO}},\cal{L}_{\text{KTO}},\ldots\}$ depending on the dataset $\mathcal{D}_{r}$
@@
5:  // Safety alignment
6:  Choose the loss function $\mathcal{L}_{g}$ within $\{\mathcal{L}_{\text{DPO}},\mathcal{L}_{\text{IPO}},\cal{L}_{\text{KTO}},\ldots\}$ depending on the dataset $\mathcal{D}_{g}$
@@
Theorem 1 (Relations between $\pi^{\star}_{r^{\star}}$ and $\pi^{\star}$).
The optimal policy of (10) is represented as
Remark 1 (Importance of Reverse KL in (10)).
Remark 2 (Commutative law).
Since the commutative law holds, alignment does not have to be conducted in the order reward to safety.
For the proof, see Appendix C. Intuitively, Theorem 1 states that we do not have to align a policy for multiple metrics simultaneously and thus we can sequentially align the policy stepwise. Specifically, (12) means that the optimal policy $\pi^{\star}$ is identical to the one obtained by realignment of $\pi^{\star}_{r^{\star}}$ for the safety function $g^{\star}$ with a parameter $\beta/\lambda^{\star}$. Thus, (12) justifies realigning the reward-aligned model $\pi^{\star}_{r^{\star}}$ with respect to the safety function $g^{\star}$. After a simple mathematical transformation, we then have
$g^{\star}(x,y)=\frac{\beta}{\lambda^{\star}}\log\frac{\pi^{\star}(y\mid x)}{\pi^{\star}_{r^{\star}}(y\mid x)}+\frac{\beta}{\lambda^{\star}}\log Y(x).$ (13)
Because the second term of the RHS does not depend on $y$ and $\pi$, for all $x\in\mathcal{X}$, we have
[…]
Observe that, the loss function slightly changed from $(\pi_{\mathrm{ref}},\beta)$ in (6) to $(\pi^{\star}_{r^{\star}},\beta/\lambda^{\star})$ in the above equation. Such modification of the loss function is valid with other algorithms such as IPO or KTO.
4.2 Advantages of SACPO
By taking a stepwise approach, we can enjoy practical benefits. Let us highlight three advantages of SACPO. The first advantage of SACPO is its flexibility of alignment algorithms (e.g., DPO or KTO) and datasets. In practice, depending on the metric, appropriate human feedback should be different (e.g., paired vs. unpaired); hence, the proper alignment algorithm or its parameters should also be different. Our proposed algorithm takes a stepwise approach to the constrained LLMs alignment, which allows us to use different algorithms, parameters, or datasets for each metric. Second, we can evaluate the resulting LM policy after each alignment with respect to the target metric, which enables us to prevent starting over the alignment from the beginning. Finally, our SACPO justifies us to realign pre-aligned LLMs with our desired metric. This is a practically desirable property because we now have easy access to high-quality, open-source LLMs that have been aligned.
@@ 5 Theoretical Results
This section provides theoretical results regarding our SACPO algorithm. Specifically, we provide upper bounds on the near-optimality and safety constraint violation. While SACPO does not explicitly estimate the reward and safety, (4) and (12) tell us that the policies are secretly reward or safety models. Hence, we first analyze the uncertainty of the estimated reward and safety functions and then derive the bounds regarding the performance of the policy obtained by SACPO. As a key notion behind our theoretical analysis, let us first introduce the uncertainty quantifier defined as follows.
@@
In the above definition, $f$ represents $r$, $g$, or their weighted summation. In RLHF pipelines, the reward model is usually initialized from the SFT model adding a linear layer on top of the final transformer layer to generate an estimated reward value. Hence, to obtain the $\delta$-uncertainty quantifier, we adopt the linear realizability assumption, which has been widely adopted in related fields such as linear bandit [23] or linear MDP [20]. Recently, Xiong et al. [43] provided theoretical analysis for RLHF and DPO under preference data and linear realizability assumptions. We extend their theoretical analysis from unconstrained to constrained settings and from preference to a more general dataset.
@@
The reward and safety functions are parameterized by $\widehat{r}(x,y)=\langle w_{r},\phi(x,y)\rangle$ and $\widehat{g}(x,y)=\langle w_{g},\phi(x,y)\rangle$ for a shared feature mapping function $\phi:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}^{d}$. In addition, the true reward and safety functions satisfy $r^{\star}(x,y)=\langle w_{r}^{\star},\phi(x,y)\rangle$ and $g(x,y)=\langle w_{g}^{\star},\phi(x,y)\rangle$ for some $w_{r}^{\star}\in\mathbb{R}^{d}$ and $w_{g}^{\star}\in\mathbb{R}^{d}$. For regularization, we have $\|\phi(x,y)\|\leq 1$ for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$ and $\max\{\|w_{r}\|,\|w_{g}\|\}\leq B$.
@@
With a dataset $\mathcal{D}$, define the covariance matrix estimation $\Sigma_{\mathcal{D}}\coloneqq\kappa\mathbb{I}+\sum_{(x,y_{1},y_{2})\in\mathcal{D}}\big{(}\phi(x,y_{1})-\phi(x,y_{2})\big{)}\big{(}\phi(x,y_{1})-\phi(x,y_{2})\big{)}^{\top}$ for paired dataset and $\Sigma_{\mathcal{D}}\coloneqq\kappa\mathbb{I}+\sum_{(x,y)\in\mathcal{D}}\phi(x,y)\phi(x,y)^{\top}$ for unpaired dataset, where $\kappa\in\mathbb{R}_{+}$ is a fixed positive value and $\mathbb{I}\in\mathbb{R}^{d\times d}$ is the identity matrix. Also, define, $\mathcal{U}_{\mathcal{D}}(x,y)\coloneqq\|\phi(x,y)\|_{\Sigma^{-1}_{\mathcal{D}}}$, where $\|\texttt{x}\|_{A}\coloneqq\sqrt{\texttt{x}^{\top}A\texttt{x}}$ is the matrix Mahalanobis seminorm. For paired dataset, with $\gamma\coloneqq 2+e^{B}+e^{-B}$, set $\alpha=\mathcal{O}\left(\sqrt{\gamma^{2}(d+\log(1/\delta))+\kappa B^{2}}\right)$. For unpaired dataset, set $\alpha=B\Bigl{(}1+\sqrt{\log(2/\delta)/2}\Bigr{)}$. Then, MLEs for reward and safety functions (i.e., $\widehat{r}$ and $\widehat{g}$) respectively satisfy
@@
Let $\widehat{\pi}$ denote the optimal policy induced by $\widehat{h}(x,y)\coloneqq\widehat{r}(x,y)+\widehat{\lambda}\widehat{g}(x,y)$; that is $\widehat{\pi}(y\mid x)=\frac{1}{Z_{\widehat{h}}(x;\pi_{\mathrm{ref}})}\pi_{\mathrm{ref}}(y\mid x)\exp(\frac{1}{\beta}(\widehat{h}(x,y)))$. Then, the following inequality holds:
@@
The second term appears hard to handle due to $\mathbb{E}_{\widehat{\pi}}[\cdot]$ but can be upper-bounded since $\frac{\widehat{\pi}(y\mid x)}{\pi^{\star}(y\mid x)}=\frac{Z_{h^{\star}}(x;\pi_{\mathrm{ref}})}{Z_{\widehat{h}}(x;\pi_{\mathrm{ref}})}\exp\left(\frac{\widehat{h}(x,y)-h^{\star}(x,y)}{\beta}\right)$ holds by definition. Because each term is upper-bounded by leveraging the uncertainty quantifiers, we can also obtain the total upper bound by adding them together.
Theorems 2 and 3 suggest two important things. First, regarding both optimality and safety constraint violation, the performance decay compared to $\pi^{\star}$ is exponential to
$\widehat{\Gamma}_{\mathcal{D}}(x,y,\lambda^{\star})=\alpha\,(\mathcal{U}_{\mathcal{D}_{r}}(x,y)+\lambda^{\star}\,\mathcal{U}_{\mathcal{D}_{g}}(x,y))+|\lambda^{\star}-\widehat{\lambda}|B.$
More concretely, to achieve high reward and safety performance, we must accurately estimate the reward and safety functions and obtain a Lagrangian multiplier of high quality. Second, the expectation is taken over $\pi^{\star}$ (i.e., $\mathbb{E}_{\pi^{\star}}$), which means that we need to collect high-quality data such that the optimal policy $\pi^{\star}$ is likely to generate.
@@ 6 Practical Implementation
An unresolved issue remains about SACPO, namely how to optimize $\lambda$. In the context of CMDP settings, since $L(\pi,\lambda,\beta)$ is linear in $\lambda$, primal-dual methods are popular for optimizing an agent’s policy and Lagrangian multiplier [13, 22]. In this process, online convex optimization approaches (e.g., [49]) are often used for optimizing $\lambda$ while evaluating the reward and safety performance of the current policy during training.
@@
As shown in Figure 1, after obtaining a reward-aligned policy $\pi_{r}$, SACPO realigns it with the safety metric $g$. In this safety-realignment phase, we set $\lambda$ a conservatively large scalar $\lambda=\bar{\lambda}$ so that $\bar{\lambda}>\lambda^{\star}$. We now own two LM policies: a reward-aligned policy $\pi_{r}$ (this can be regarded as a policy with $\lambda=0$) and a conservatively safety-realigned policy $\pi_{r+\bar{\lambda}g}$. Under the assumption that $\lambda^{\star}$ is between 00 and $\bar{\lambda}$, we aim to find $\lambda^{\star}$ without optimizing new LM policies. Specifically, we merge $\pi_{r}$ and $\pi_{r+\bar{\lambda}g}$ by simply averaging their weights as in Wortsman et al. [42] with a mixing ratio of $q:1-q$ for a given scalar $q\in\mathbb{R}_{+}$ ($0\leq q\leq 1$).

### v2

@@ 4 Stepwise Alignment for Constrained Policy Optimization
We propose an algorithm called SACPO to solve the constrained policy optimization problem (9), which is outlined in Algorithm 1. SACPO takes a stepwise approach for reward and safety alignments of LLMs; that is, an LLM is first aligned for reward and then for safety (or vice versa). This operation is backed by theory, and the resulting optimal LM policy is guaranteed to be identical to the one aligned with reward and safety metrics simultaneously. In addition, by taking the stepwise approach, we can enjoy several practical advantages of enhanced flexibility regarding algorithms and datasets. Though we formulate a problem with a single safety function for simplicity, the arguments in this section are valid in the case of multiple safety functions. For more details, see Appendix B.2 and B.3.
SACPO uses a standard Lagrangian [8] defined as $L(\pi,\lambda,\beta)\coloneqq R(\pi,\beta)+\lambda(G(\pi)-b)$, where $\pi\in\Pi$ is the primal variable and $\lambda\in\mathbb{R}_{+}$ is a dual variable or the Lagrangian multiplier. Note that, for any dual variable $\lambda\in\mathbb{R}_{+}$, by using a composite function $r^{\star}+\lambda g^{\star}$, we can convert the original constrained policy optimization problem (9) into the following max-min problem:
$\max_{\pi}\min_{\lambda\geq 0}\ L(\pi,\lambda,\beta)=\mathbb{E}_{\rho,\pi}\bigl{[}r^{\star}(x,y)+\lambda g^{\star}(x,y)\bigr{]}-\beta\mathbb{D}_{\mathrm{KL}}[\,\pi_{\theta}(y\mid x)\,\|\,\pi_{\mathrm{ref}}(y\mid x)\,]-\lambda b.$ (10)
@@
Practically, it is not hard to obtain such a conservative policy $\overline{\pi}$. If the usefulness (i.e., reward $r$) can be ignored, it is easy to acquire policies that refuse to generate potentially unsafe answers and output safe answers conservatively. Based on Assumption 1, we present the following two lemmas.
@@
For the proofs, see Appendix C. Our problem setting is a special case of those in typical constrained Markov decision process (CMDP, [4]) literature. Thus, Lemma 1 follows from Theorem 3 in Paternain et al. [34], and Lemma 2 follows from Lemma 1 in Ding et al. [16].
@@ 4.1 Optimal Policy Can be Directly Obtained from Reward-aligned Policy
Theorem 1 (Relation between $\pi^{\star}_{r^{\star}}$ and $\pi^{\star}$).
The optimal policy of (9) is represented as
Remark 1 (Importance of reverse KL in (3) and (9)).
Remark 2 (Commutative law).
Since the commutative law holds, alignment does not have to be conducted in the order from reward to safety.
For the proof, see Appendix D. Intuitively, Theorem 1 states that we do not have to align a policy for multiple metrics simultaneously and thus we can sequentially align the policy stepwise. Specifically, (11) means that the optimal policy $\pi^{\star}$ is identical to the one obtained by realignment of $\pi^{\star}_{r^{\star}}$ for the safety function $g^{\star}$ with a parameter $\beta/\lambda^{\star}$. Thus, (11) justifies realigning the reward-aligned model $\pi^{\star}_{r^{\star}}$ with respect to the safety function $g^{\star}$. After a simple mathematical transformation of (11) and the fact that $\log Y(x)$ neither depend on $y$ nor $\pi$ for all $x\in\mathcal{X}$, we then have
[…]
The loss function slightly changed from $(\pi_{\mathrm{ref}},\beta)$ in (6) to $(\pi^{\star}_{r^{\star}},\beta/\lambda^{\star})$. Such modification of the loss function is valid with other algorithms that explicitly use (4) for the deviation (e.g., IPO, KTO).
Advantages of SACPO. By taking a stepwise approach, we can enjoy practical benefits. Let us highlight three major advantages of SACPO. The first advantage is the flexibility of alignment algorithms (e.g., DPO or KTO) and datasets. In practice, depending on the metric, appropriate human feedback should be different (e.g., paired vs. unpaired). SACPO takes a stepwise approach, which allows us to use different algorithms, parameters, or datasets for each metric. Second, we can evaluate the resulting LM policy after each alignment regarding the target metric. This process enables us to prevent starting over the alignment from the beginning. Finally, SACPO justifies us to realign pre-aligned LLMs with our desired metric. This property is practically desirable because we now have easy access to high-quality, open-source LLMs that have been already aligned.
@@
3:  Choose the loss function $\mathcal{L}_{r}$ within $\{\mathcal{L}_{\text{DPO}},\cal{L}_{\text{KTO}},\ldots\}$ depending on the dataset $\mathcal{D}_{r}$
@@
5:  // Safety realignment
6:  Choose the loss function $\mathcal{L}_{g}$ within $\{\mathcal{L}_{\text{DPO}},\cal{L}_{\text{KTO}},\ldots\}$ depending on the dataset $\mathcal{D}_{g}$
@@ 5 Theoretical Results
This section provides theoretical results. Specifically, we provide the upper bounds on the optimality and safety constraint violation of the policy obtained by SACPO. While SACPO does not explicitly estimate the reward and safety, (4) and (11) tell us that the policies are secretly reward or safety models. Hence, we first analyze the uncertainty of the estimated reward and safety functions and then derive the bounds on the performance of the policy trained via SACPO. As a key notion in our theoretical analysis, let us define an uncertainty quantifier as follows.
@@
Note that $f$ represents $r$, $g$, or their weighted summation. In RLHF pipelines, the reward model is usually initialized from the SFT model by adding a linear layer on top of the final transformer layer to generate an estimated reward value. Recently, Xiong et al. [50] have provided theoretical analysis for RLHF and DPO under preference data and linear realizability assumptions. We extend their theory from unconstrained to constrained settings and from preference to a more general dataset.
@@
The reward and safety functions are parameterized by $\widehat{r}(x,y)=\langle w_{r},\phi(x,y)\rangle$ and $\widehat{g}(x,y)=\langle w_{g},\phi(x,y)\rangle$ for a shared feature mapping function $\phi:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}^{d}$. In addition, the true reward and safety functions satisfy $r^{\star}(x,y)=\langle w_{r}^{\star},\phi(x,y)\rangle$ and $g(x,y)=\langle w_{g}^{\star},\phi(x,y)\rangle$ for some $w_{r}^{\star},w_{g}^{\star}\in\mathbb{R}^{d}$. For regularization, we additionally assume $\|\phi(x,y)\|\leq 1$ for any $(x,y)\in\mathcal{X}\times\mathcal{Y}$ and $\max\{\|w_{r}\|,\|w_{g}\|\}\leq B$.
@@
With a dataset $\mathcal{D}$, define the covariance matrix estimation $\Sigma_{\mathcal{D}}\coloneqq\kappa\mathbb{I}+\sum_{(x,y_{1},y_{2})\in\mathcal{D}}\big{(}\phi(x,y_{1})-\phi(x,y_{2})\big{)}\big{(}\phi(x,y_{1})-\phi(x,y_{2})\big{)}^{\top}$ for paired dataset and $\Sigma_{\mathcal{D}}\coloneqq\kappa\mathbb{I}+\sum_{(x,y)\in\mathcal{D}}\phi(x,y)\phi(x,y)^{\top}$ for unpaired dataset, where $\kappa\in\mathbb{R}_{+}$ is a fixed positive value and $\mathbb{I}\in\mathbb{R}^{d\times d}$ is the identity matrix. Also, define, $\mathcal{U}_{\mathcal{D}}(x,y)\coloneqq\|\phi(x,y)\|_{\Sigma^{-1}_{\mathcal{D}}}$, where $\|\texttt{x}\|_{A}\coloneqq\sqrt{\texttt{x}^{\top}A\texttt{x}}$ is the matrix Mahalanobis seminorm. For paired dataset, with $\gamma\coloneqq 2+e^{B}+e^{-B}$, set $\alpha=\mathcal{O}(\sqrt{\gamma^{2}(d+\log(1/\delta))+\kappa B^{2}})$. For unpaired dataset, set $\alpha=B(1+\sqrt{\log(2/\delta)/2})$. Then, MLEs for reward and safety functions (i.e., $\widehat{r}$ and $\widehat{g}$) respectively satisfy
@@
$\widehat{\Gamma}_{\mathcal{D}}(x,y,c)$ $\coloneqq\alpha\left(\,\mathcal{U}_{\mathcal{D}_{r}}(x,y)+c\,\mathcal{U}_{\mathcal{D}_{g}}(x,y)\right)+|\,c-\widehat{\lambda}\,|B,$ (12)
@@
Let $\widehat{\pi}$ denote the optimal policy induced by $\widehat{h}(x,y)\coloneqq\widehat{r}(x,y)+\widehat{\lambda}\widehat{g}(x,y)$; that is $\widehat{\pi}(y\mid x)=\frac{1}{Z_{\widehat{h}}(x;\pi_{\mathrm{ref}})}\pi_{\mathrm{ref}}(y\mid x)\exp\big{(}\frac{1}{\beta}(\widehat{h}(x,y))\big{)}$. Then, the following inequality holds:
@@
$J_{f}(\pi^{\star})-J_{f}(\widehat{\pi})=\mathbb{E}_{\rho}\left[{\mathbb{E}_{\pi^{\star}}\bigl{[}f(x,y)-h^{\star}(x,y)\bigr{]}}+{\mathbb{E}_{\widehat{\pi}}\bigl{[}\,\widehat{h}(x,y)-f(x,y)\bigr{]}}+\beta\log\frac{Z_{h^{\star}}(x;\pi_{\mathrm{ref}})}{Z_{\widehat{h}}(x;\pi_{\mathrm{ref}})}\right].$
The second term on the right-hand side appears hard to handle due to $\mathbb{E}_{\widehat{\pi}}[\cdot]$ but is upper-bounded since $\frac{\widehat{\pi}(y\mid x)}{\pi^{\star}(y\mid x)}=\frac{Z_{h^{\star}}(x;\pi_{\mathrm{ref}})}{Z_{\widehat{h}}(x;\pi_{\mathrm{ref}})}\exp\left(\frac{\widehat{h}(x,y)-h^{\star}(x,y)}{\beta}\right)$ holds by definition. Because each term is upper-bounded by leveraging the uncertainty quantifiers, we obtain the total upper bound by adding them together.
Theorems 2 and 3 suggest that, regarding both optimality and safety violation, the performance degradation compared to $\pi^{\star}$ is exponential to $\frac{2}{\beta}\,\widehat{\Gamma}_{\mathcal{D}}(x,y,\lambda^{\star})=\frac{2}{\beta}\big{(}\alpha\,\bigl{(}\mathcal{U}_{\mathcal{D}_{r}}(x,y)+\lambda^{\star}\,\mathcal{U}_{\mathcal{D}_{g}}(x,y)\bigr{)}+|\lambda^{\star}-\widehat{\lambda}|B\big{)}$. Our theoretical results imply that accurate estimation of the reward and safety functions and a high-quality Lagrangian multiplier are required to achieve high reward and safety performance.
@@ 6 Practical Implementation
An unresolved issue remains about SACPO, namely how to optimize $\lambda$. In the typical CMDP settings, since $L(\pi,\lambda,\beta)$ is linear in $\lambda$, primal-dual methods are popular for optimizing $\pi$ and $\lambda$ [16, 26]. In this process, online convex optimization approaches (e.g., [56]) are often used for optimizing $\lambda$ while evaluating the reward and safety performance of the current policy during training.
@@
We now introduce a practical variant of SACPO  called P-SACPO. After obtaining a reward-aligned policy $\pi_{r}$, P-SACPO realigns it with the safety metric $g$ while setting $\lambda$ as a conservatively large scalar $\lambda=\bar{\lambda}$ such that $\bar{\lambda}>\lambda^{\star}$. We now own two LM policies: a reward-aligned policy $\pi_{r}$ (this can be regarded as an LM policy with $\lambda=0$) and a conservatively safety-realigned policy $\pi_{r+\bar{\lambda}g}$. Under the assumption that $\lambda^{\star}$ is between 00 and $\bar{\lambda}$, P-SACPO aims to find $\lambda^{\star}$ without optimizing new LM policies. Specifically, we merge $\pi_{r+\bar{\lambda}g}$ and $\pi_{r}$ by simply averaging their weights as in Wortsman et al. [49] with a mixing ratio of $q:1-q$ for a scalar $q\in\mathbb{R}_{+}$ ($0\leq q\leq 1$). It is known that such a simple weight-averaging works well in the case of the same base model [18, 3]. All the models obtained by SACPO derive from the same SFT model. Therefore, SACPO is particularly compatible with model merging, and P-SACPO empirically performs well as evidenced by our experiments in Section 7.